OpenStudy (anonymous):

Find the x and y intercepts of the equation algebraically. 4x-3y=10

OpenStudy (anonymous):

The x intercept will look be (0,y) and the y intercept will be (x,0), so: 4(0) -3y =10, y = -10/3 For the x intercept, 4x -3(0) = 10 x = 10/4. Y intercept = (0,-10/3) X intercept = (10/4,0)

OpenStudy (anonymous):

Ok, thank you! I wanted to be sure that I did it right! :)